OpenStudy (anonymous):

If an object is falling with a position function P(t)=-16t^2+2000 how do i A) find the average volocity on the interval [4,5]? B)show that at t=4.5, the velocity is the same as the average of the interval.

OpenStudy (phi):

For (A), plug in 4 for t in your equation to find the position at time 4. Then do the same for t=5. The difference is how far it moved in 1 second, or its average velocity. for (B), plug in t= 4.5 into the equation.

OpenStudy (anonymous):

\[f_{ave}=1/(5-4)\int\limits_{4}^{5}P'(t)dt\] =[P(t)]4,5 =P(5)-P(4)